# Unveiling Impact of Frequency Components on Membership Inference Attacks for Diffusion Models

### v1

@@ Appendix C Proof of Theorem 1
For a given sample $x_{i}$, the membership score can be obtained according to the general paradigm we have defined:
$s=||x_{i,t}-x_{i,t}^{target}||.$ (C.2)
Following [42], let the error conditional probability density functions $f(s|m=1)$ and $f(s|m=0)$:
$f(s|m=1)=\frac{1}{\sqrt{2\pi\sigma_{M}}}e^{-s^{2}/(2\sigma_{M}^{2})},$ (C.3)
$f(s|m=0)=\frac{1}{\sqrt{2\pi\sigma_{H}}}e^{-s^{2}/(2\sigma_{H}^{2})}.$ (C.4)
After knowing $f(s|m=0)$ and $f(s|m=1)$, the attacker selects the maximum likelihood of $f(s|m)$ as the output. This is based on determining the probability density function to decide whether the data point $x$ comes from the member data $(m=1)$ or the hold-out data $(m=0)$. Given $f(s|m=0)=f(s|m=1)$, the formula is equivalent to:
$\frac{1}{\sqrt{2\pi\sigma_{M}}}e^{-s^{2}/(2\sigma_{M}^{2})}=\frac{1}{\sqrt{2\pi\sigma_{H}}}e^{-s^{2}/(2\sigma_{H}^{2})}.$ (C.5)
By the properties of logarithms, we obtain:
$\ln 1-\ln(\sqrt{2\pi}\sigma_{M})+\ln\left(e^{-s^{2}/(2\sigma_{M}^{2})}\right)=\ln 1-\ln(\sqrt{2\pi}\sigma_{H})+\ln\left(e^{-s^{2}/(2\sigma_{H}^{2})}\right),$ (C.6)
@@
$-\frac{s^{2}}{2\sigma_{M}^{2}}-\ln(\sigma_{M})=-\frac{s^{2}}{2\sigma_{H}^{2}}-\ln(\sigma_{H}).$ (C.7)
Let ±$\pm s_{eq}$ be the points at which these two probability density functions are equal:
$s_{eq}=\sigma_{H}\sqrt{\frac{2\ln(\sigma_{H}/\sigma_{M})}{(\sigma_{H}/\sigma_{M})^{2}-1}}.$ (C.8)
If $\sigma_{M}<\sigma_{H}$, $f(s|m=1)>f(s|m=0)$ if and only if $|s|<s_{eq}$. The membership advantage, as defined in Definition 1, can be rewritten as:
$Adv^{M}(\mathcal{A})=\Pr[|s|<s_{eq}|m=1]-\Pr[|s|<s_{eq}|m=0].$ (C.9)
According to the relationship between the error function and the standard normal distribution, Eq. C.9 can be converted to:
$Adv^{M}(\mathcal{A})=erf\left(\frac{s_{eq}}{\sqrt{2\sigma_{M}}}\right)-erf\left(\frac{s_{eq}}{\sqrt{2\sigma_{H}}}\right).$ (C.10)
Substituting $s_{eq}$ into the Eq. C.10 yields:
$Adv^{M}(\mathcal{A})=erf\left(\frac{\sigma_{H}}{\sigma_{M}}\sqrt{\frac{\ln(\sigma_{H}/\sigma_{M})}{(\sigma_{H}/\sigma_{M})^{2}-1}}\right)-erf\left(\sqrt{\frac{\ln(\sigma_{H}/\sigma_{M})}{(\sigma_{H}/\sigma_{M})^{2}-1}}\right).$ (C.11)
Therefore, Theorem 1 is proved. ∎
Appendix D Proof of Proposition 1
Proposition 1.
Denote the original standard deviations of membership scores in member and hold-out data as $\sigma_{M}$ and $\sigma_{H}$, and the standard deviations after removing the high-frequency components are $\sigma_{M}^{\prime}$ and $\sigma_{H}^{\prime}$. The standard deviation of membership scores in the high-frequency components is $h_{M}/h_{H}$, and the low-frequency components is $l_{M}/l_{H}$ in member and hold-out data. Let $l_{H}-l_{M}=\Delta$, $h_{M}=k\cdot h_{H}$ with $k>0$. If $k^{2}>1+\frac{2\Delta}{h_{H}^{2}}(l_{M}+2\Delta-\sqrt{(l_{M}+2\Delta)^{2}+h_{H}^{2}})$, we have:
$\sigma_{H}^{\prime}/\sigma_{M}^{\prime}>\sigma_{H}/\sigma_{M}.$ (D.1)
Proof.
We follow the premise of Theorem 1, the standard deviation before and after filtering high-frequency information satisfies $\sigma_{H}>\sigma_{M}$ and $\sigma_{H}^{\prime}>\sigma_{M}^{\prime}$. $\frac{\sigma_{H}^{\prime}}{\sigma_{M}^{\prime}}>\frac{\sigma_{H}}{\sigma_{M}}$ is equivalent to:
$\sigma_{H}^{\prime}\sigma_{M}>\sigma_{H}\sigma_{M}^{\prime}.$ (D.2)
Let $\sigma_{H}=\sigma_{H}^{\prime}+\Delta_{H}$ and $\sigma_{M}=\sigma_{M}^{\prime}+\Delta_{M}$. Substituting these into Eq. D.2, we obtain:
$\sigma_{H}^{\prime}\sigma_{M}^{\prime}+\sigma_{H}^{\prime}\Delta_{M}>\sigma_{H}^{\prime}\sigma_{M}^{\prime}+\Delta_{H}\sigma_{M}^{\prime}.$ (D.3)
Since $\sigma_{H}^{\prime}>\sigma_{M}^{\prime}>0$, Eq. D.3 holds constantly if and only if:
$\Delta_{M}-\Delta_{H}>0.$ (D.4)
$\Delta_{M}=\sqrt{l_{M}^{2}+h_{M}^{2}+cov(l_{M},h_{M})}-\sqrt{l_{M}^{2}},\Delta_{H}=\sqrt{l_{H}^{2}+h_{H}^{2}+cov(l_{H},h_{H})}-\sqrt{l_{H}^{2}}.$ (D.5)
Eq. D.4 can be expressed as:
$\sqrt{l_{M}^{2}+h_{M}^{2}+cov(l_{M},h_{M})}-\sqrt{l_{H}^{2}+h_{H}^{2}+cov(l_{H},h_{H})}>l_{M}-l_{H}.$ (D.6)
$\sqrt{l_{M}^{2}+h_{M}^{2}}>\sqrt{(l_{M}+\Delta)^{2}+h_{H}^{2}}-\Delta.$ (D.7)
Square both sides of Eq. D.7 to obtain the following:
$l_{M}^{2}+h_{H}^{2}>(l_{M}+\Delta)^{2}+h_{H}^{2}-2\Delta\sqrt{(l_{M}+\Delta)^{2}+h_{M}^{2}}+\Delta^{2}.$ (D.8)
Substitute $h_{M}=k\cdot h_{H}$ into Eq. D.8, then:
$(k^{2}-1)h_{H}^{2}>2l_{M}\Delta+2\Delta^{2}-2\Delta\sqrt{(l_{M}+\Delta)^{2}+k^{2}h_{H}^{2}}.$ (D.9)
Add $(l_{M}+\Delta)^{2}+h_{H}^{2}$ on both sides of Eq. D.9:
$(l_{M}+\Delta)^{2}+k^{2}h_{H}^{2}>2l_{M}\Delta+2\Delta^{2}-2\Delta\sqrt{(l_{M}+\Delta)^{2}+k^{2}h_{H}^{2}}+(l_{M}+\Delta)^{2}+h_{H}^{2}.$ (D.10)
Let $t=(l_{M}+\Delta)^{2}+k^{2}h_{H}^{2}$, substitute $t$ into Eq. D.10:
$t>2l_{M}\Delta+2\Delta^{2}-2\Delta\sqrt{t}+(l_{M}+\Delta)^{2}+h_{H}^{2},$ (D.11)
which is equivalent to:
$(\sqrt{t}+\Delta)^{2}>(l_{M}+2\Delta)^{2}+h_{H}^{2}.$ (D.12)
@@ Appendix D Proof of Proposition 1
$\sqrt{t}>\sqrt{(l_{M}+2\Delta)^{2}+h_{H}^{2}}-\Delta.$ (D.13)
Substitute back $t=(l_{M}+\Delta)^{2}+k^{2}h_{H}^{2}$ to Eq. D.13, we obtain:
$(l_{M}+\Delta)^{2}+k^{2}h_{H}^{2}>(l_{M}+2\Delta)^{2}+h_{H}^{2}-2\Delta\sqrt{(l_{M}+2\Delta)^{2}+h_{H}^{2}}+\Delta^{2}.$ (D.14)
@@
$k^{2}>1+\frac{2\Delta}{h_{H}^{2}}(l_{M}+2\Delta-\sqrt{(l_{M}+2\Delta)^{2}+h_{H}^{2}}).$ (D.15)
@@
Appendix E Complementary Experiments
E.1 More Frequency Domain Analysis
@@
Figure 5: PIA pixel-wise errors distribution visualization.
E.2 Detailed settings
@@
E.3 Attack for Pre-trained Stable Diffusion
@@
E.4 Membership Scores Distribution for Samples from Member and Hold-out Set.
@@
E.5 Ablation Study on DDIM
@@
E.6 Missing Text
Previously, for attacks on stable diffusion, we typically assumed access to the text used for each image during training, which is a strong requirement for attackers. In real-world scenarios, we may not be able to obtain the text, which increases the difficulty of the attack. To simulate more realistic attack conditions, we conducted attacks in scenarios without text and with text generated by BLIP [21]. As shown in Tab. 12, when attacking without text or with the text generated by BLIP, although the performance of the baselines has declined noticeably, the filter still shows satisfactory results. This demonstrates that our method still exhibits good performance under more stringent attack conditions.
Method MS-COCO Flickr
No Text BLIP Text No Text BLIP Text
ASR AUC ASR AUC ASR AUC ASR AUC
Naive 61.19 64.45 75.80 81.43 69.74 73.80 75.49 80.99
Naive+F 69.80 74.56 87.72 91.51 81.75 88.23 87.75 94.77
SecMI 69.60 74.77 81.85 87.99 75.00 79.22 77.99 84.00
SecMI+F 70.70 76.70 82.92 88.76 75.67 80.26 78.95 84.87
PIA 53.10 51.99 61.42 65.69 57.24 58.75 66.00 66.84
PIA+F 54.70 53.18 64.94 70.81 59.61 59.25 69.00 73.29
Avg+ +3.77 +4.41 +5.50 +5.32 +5.02 +5.32 +5.41 +7.03
E.7 DDIM without data augmentation
As shown in Tab. 13, we tested the DDIM trained without data augmentation. The results indicate that our method still demonstrates significant effects. It is worth noting that on the CIFAR-100 dataset, existing attacks have already achieved outstanding performance in terms of ASR and AUC metrics; therefore, our method shows minor improvements in these two metrics. However, there is an average improvement of 18.58% in the TPR@1% FPR metric.
Table 13: After removing data augmentation from training, the attack performance in DDIM.
Method STL10-U CIFAR-100 Tiny-IN
ASR AUC TPR@1%FPR ASR AUC TPR@1%FPR ASR AUC TPR@1%FPR
Naive 81.82 88.75 10.92 94.02 97.98 47.65 80.00 87.61 11.75
Naive+F 85.22 91.89 13.84 96.46 99.01 66.98 86.07 93.09 21.40
SecMI 85.34 91.95 18.20 95.70 98.93 71.19 83.77 90.96 20.63
SecMI+F 91.93 96.82 37.13 98.83 99.85 98.16 93.56 98.04 61.24
PIA 88.88 94.98 29.63 97.54 99.54 90.48 88.06 94.49 33.88
PIA+F 93.14 97.54 45.68 99.13 99.91 99.91 94.41 98.52 72.54
Avg+ +4.75 +3.52 +12.63 +2.39 +0.77 +18.58 +7.40 +5.53 +29.64
E.8 ROC curves Visualization
@@ E.9 Compute Overhead
Appendix F Ethics Statements

### v2

@@ Appendix C Proof of Proposition 1
The standard deviation before and after filtering high-frequency information satisfies $\sigma_{H}>\sigma_{M}$ and $\sigma_{H}^{\prime}>\sigma_{M}^{\prime}$. $\frac{\sigma_{H}^{\prime}}{\sigma_{M}^{\prime}}>\frac{\sigma_{H}}{\sigma_{M}}$ is equivalent to:
$\sigma_{H}^{\prime}\sigma_{M}>\sigma_{H}\sigma_{M}^{\prime}.$ (C.2)
Let $\sigma_{H}=\sigma_{H}^{\prime}+\Delta_{H}$ and $\sigma_{M}=\sigma_{M}^{\prime}+\Delta_{M}$. Substituting these into Eq. C.2, we obtain:
$\sigma_{H}^{\prime}\sigma_{M}^{\prime}+\sigma_{H}^{\prime}\Delta_{M}>\sigma_{H}^{\prime}\sigma_{M}^{\prime}+\Delta_{H}\sigma_{M}^{\prime}.$ (C.3)
Since $\sigma_{H}^{\prime}>\sigma_{M}^{\prime}>0$, Eq. C.3 holds constantly if and only if:
$\Delta_{M}-\Delta_{H}>0.$ (C.4)
$\Delta_{M}=\sqrt{l_{M}^{2}+h_{M}^{2}+cov(l_{M},h_{M})}-\sqrt{l_{M}^{2}},\Delta_{H}=\sqrt{l_{H}^{2}+h_{H}^{2}+cov(l_{H},h_{H})}-\sqrt{l_{H}^{2}}.$ (C.5)
Eq. C.4 can be expressed as:
$\sqrt{l_{M}^{2}+h_{M}^{2}+cov(l_{M},h_{M})}-\sqrt{l_{H}^{2}+h_{H}^{2}+cov(l_{H},h_{H})}>l_{M}-l_{H}.$ (C.6)
$\sqrt{l_{M}^{2}+h_{M}^{2}}>\sqrt{(l_{M}+\Delta)^{2}+h_{H}^{2}}-\Delta.$ (C.7)
Square both sides of Eq. C.7 to obtain the following:
$l_{M}^{2}+h_{H}^{2}>(l_{M}+\Delta)^{2}+h_{H}^{2}-2\Delta\sqrt{(l_{M}+\Delta)^{2}+h_{M}^{2}}+\Delta^{2}.$ (C.8)
Substitute $h_{M}=k\cdot h_{H}$ into Eq. C.8, then:
$(k^{2}-1)h_{H}^{2}>2l_{M}\Delta+2\Delta^{2}-2\Delta\sqrt{(l_{M}+\Delta)^{2}+k^{2}h_{H}^{2}}.$ (C.9)
Add $(l_{M}+\Delta)^{2}+h_{H}^{2}$ on both sides of Eq. C.9:
$(l_{M}+\Delta)^{2}+k^{2}h_{H}^{2}>2l_{M}\Delta+2\Delta^{2}-2\Delta\sqrt{(l_{M}+\Delta)^{2}+k^{2}h_{H}^{2}}+(l_{M}+\Delta)^{2}+h_{H}^{2}.$ (C.10)
Let $t=(l_{M}+\Delta)^{2}+k^{2}h_{H}^{2}$, substitute $t$ into Eq. C.10:
$t>2l_{M}\Delta+2\Delta^{2}-2\Delta\sqrt{t}+(l_{M}+\Delta)^{2}+h_{H}^{2},$ (C.11)
@@
$(\sqrt{t}+\Delta)^{2}>(l_{M}+2\Delta)^{2}+h_{H}^{2}.$ (C.12)
@@
$\sqrt{t}>\sqrt{(l_{M}+2\Delta)^{2}+h_{H}^{2}}-\Delta.$ (C.13)
Substitute back $t=(l_{M}+\Delta)^{2}+k^{2}h_{H}^{2}$ to Eq. C.13, we obtain:
$(l_{M}+\Delta)^{2}+k^{2}h_{H}^{2}>(l_{M}+2\Delta)^{2}+h_{H}^{2}-2\Delta\sqrt{(l_{M}+2\Delta)^{2}+h_{H}^{2}}+\Delta^{2}.$ (C.14)
@@
$k^{2}>1+\frac{2\Delta}{h_{H}^{2}}(l_{M}+2\Delta-\sqrt{(l_{M}+2\Delta)^{2}+h_{H}^{2}}).$ (C.15)
@@
Appendix D Complementary Experiments
D.1 More Frequency Domain Analysis
@@
Figure 5: PIA pixel-wise errors distribution visualization. The areas of high error often coincide with areas of high-frequency information.
D.2 Discussion of the theoretical analysis
Constraint Conditions. We conducted experimental validation of the constraint in Proposition 1. As shown in Tab. 7, the parameter $k$ satisfies the constraint across different datasets and models, and in many cases, it even holds that $k>1$. This further demonstrates the practical feasibility and effectiveness of our theoretical results.
Method
Tiny-IN STLU-10 MS-COCO Flickr
$k^{2}$ $f$ $k^{2}$ $f$ $k^{2}$ $f$ $k^{2}$ $f$
Naive 0.988 >> 0.921 0.976 >> 0.932 1.139 >> 0.924 0.910 >> 0.892
PIA 0.976 >> 0.927 1.290 >> 0.904 1.290 >> 0.914 1.317 >> 0.909
SecMI 1.102 >> 0.920 0.959 >> 0.897 1.778 >> 0.919 1.102 >> 0.911
Normality Test. The visualization of membership score distributions in Fig. 6 shows that the score distributions for member and hold-out data exhibit an approximately symmetric unimodal characteristic, aligning with the fundamental traits of a normal distribution (centrality and symmetry). To further ascertain the conformity of the scores to a normal distribution, we conducted the Kolmogorov-Smirnov normality test on the Flickr and MS-COCO datasets. The results of this test, presented in Tab. 8, demonstrate that under a significance level of $\alpha=0.05$, the scores of baselines conform to a normal distribution, thereby affirming the validity of our hypothesis.
Table 8: The normality tests on the membership scores of the baselines. The results show that, at a significance level of $\alpha=0.05$, the baseline scores conform to a normal distribution.
Method
MS-COCO Flickr
Member Hold-out Member Hold-out
Test Statistic P-value Test Statistic P-value Test Statistic P-value Test Statistic P-value
Naive 0.0285 0.8003 0.0435 0.2921 0.0279 0.8212 0.0464 0.2248
SecMI 0.0359 0.5295 0.0428 0.3104 0.0557 0.0862 0.0446 0.1031
PIA 0.0384 0.4406 0.0327 0.6467 0.0254 0.8948 0.0449 0.2587
The Impact of Normality. By comparing the p-values from the normality tests, we observe that normality does not significantly influence our method. As illustrated in Tab. 8, the baselines exhibit varying degrees of normality across different datasets, with SecMI occasionally demonstrating weaker normality and PIA’s normality appearing somewhat inconsistent. Nevertheless, our approach is not affected by normality, as shown in Tab. 9.
Table 9: Impact of normality on our method. The results indicate that our method does not strongly rely on the normality assumption.
Method
MS-COCO Flickr
ASR AUC TPR@1%FPR ASR AUC TPR@1%FPR
SecMI 82.09 89.37 16.79 71.49 77.31 6.19
SecMI+F 91.00 95.74 27.40 80.10 85.95 21.20
Gain +8.91 +6.37 +10.61 +8.61 +8.64 +15.01
PIA 68.19 72.88 5.20 64.60 67.95 5.79
PIA+F 76.00 83.08 16.59 69.30 74.63 19.60
Gain +7.81 +10.20 +11.39 +4.70 +6.68 +13.81
D.3 Detailed settings
@@
D.4 Attack for Pre-trained Stable Diffusion
@@
D.5 The Generality of High-Frequency Filtering
The baseline methods (Naive, SecMI, and PIA) rely on the reconstruction errors of images when calculating membership scores, and our method significantly improves their performance. In fact, any attack that relies on the model’s reconstruction capability can consider employing our method. We also add experimental results for CLID (Zhai et al., 2024), which is based on likelihood estimation. Testing under the assumption of weak overfitting as outlined in Sec. 5.5. As shown in Tab. 12, it demonstrates that our method remains effective and shows significant performance improvements across different datasets. Overall, any methodology leveraging the model’s capabilities for image processing can benefit from the frequency analysis and the techniques for mitigating high-frequency deficiency presented in this paper, as this flaw is inherently present in diffusion models.
Table 12: Attack performance of CLID (Zhai et al., 2024) under weak overfitting conditions.
Method
Pokémon MS-COCO Flickr
ASR AUC TPR@1%FPR ASR AUC TPR@1%FPR ASR AUC TPR@1%FPR
CLID 87.71 91.14 54.20 89.52 93.32 59.75 88.76 92.00 56.00
CLID+F 89.75 93.73 60.25 92.10 95.41 66.00 91.20 94.53 64.00
D.6 Membership Scores Distribution for Samples from Member and Hold-out Set.
@@ D.5 The Generality of High-Frequency Filtering
D.7 Missing Text
Previously, for attacks on stable diffusion, we typically assumed access to the text used for each image during training, which is a strong requirement for attackers. In real-world scenarios, we may not be able to obtain the text, which increases the difficulty of the attack. To simulate more realistic attack conditions, we conducted attacks in scenarios without text and with text generated by BLIP (Li et al., 2022). As shown in Tab. 14, when attacking without text or with the text generated by BLIP, although the performance of the baselines has declined noticeably, the filter still shows satisfactory results. This demonstrates that our method still exhibits good performance under more stringent attack conditions.
Method
MS-COCO Flickr
No Text BLIP Text No Text BLIP Text
ASR AUC ASR AUC ASR AUC ASR AUC
Naive 61.19 64.45 75.80 81.43 69.74 73.80 75.49 80.99
Naive+F 69.80 74.56 87.72 91.51 81.75 88.23 87.75 94.77
SecMI 69.60 74.77 81.85 87.99 75.00 79.22 77.99 84.00
SecMI+F 70.70 76.70 82.92 88.76 75.67 80.26 78.95 84.87
PIA 53.10 51.99 61.42 65.69 57.24 58.75 66.00 66.84
PIA+F 54.70 53.18 64.94 70.81 59.61 59.25 69.00 73.29
Avg+ +3.77 +4.41 +5.50 +5.32 +5.02 +5.32 +5.41 +7.03
D.8 Ablation Study on DDIM
@@
D.9 ROC curves Visualization
@@ D.10 Compute Overhead
D.11 The impact of defense
We further present results on the Pokémon dataset to illustrate the impact of defense mechanisms on our method. As shown in Tab. 17 and Tab. 18, our method remains effective with the addition of defenses and provides significant improvements over the baselines.
Table 17: Attack performance ASR under the defenses. Our method remains effective and continues to deliver a clear performance improvement on ASR.
$SS_{e_{i}}$ DataAug Naive Naive+F Gain SecMI SecMI+F Gain PIA PIA+F Gain
$\times$ $\times$ 80.00 92.00 +12.00 81.08 86.00 +4.92 75.49 87.00 +11.51
$\times$ $\checkmark$ 79.50 87.88 +8.38 76.37 83.75 +7.38 72.27 80.87 +8.60
$\checkmark$ $\times$ 67.91 73.12 +5.21 77.52 84.01 +6.49 63.32 68.61 +5.29
$\checkmark$ $\checkmark$ 65.75 70.13 +4.38 74.12 79.00 +4.88 59.12 64.63 +5.51
Table 18: Attack performance TPR@1%FPR under the defenses. Our method remains effective and continues to deliver a clear performance improvement on TPR@1%FPR.
$SS_{e_{i}}$ DataAug Naive Naive+F Gain SecMI SecMI+F Gain PIA PIA+F Gain
$\times$ $\times$ 6.99 46.00 +42.01 15.30 33.00 +17.70 10.60 38.00 +27.40
$\times$ $\checkmark$ 6.49 41.25 +34.76 12.74 31.25 +18.51 7.75 39.25 +31.50
$\checkmark$ $\times$ 6.38 26.41 +20.03 12.87 25.12 +12.25 6.37 18.74 +12.37
$\checkmark$ $\checkmark$ 5.24 20.50 +15.26 11.24 22.75 +11.51 5.24 13.75 +8.51
Appendix E Ethics Statements
@@
Appendix F The Use of Large Language Models
During the preparation of this manuscript, we used a large language model (LLM) to improve the clarity and readability of the text. The use of the LLM was limited to language editing, and it did not contribute to the scientific content, analysis, or conclusions presented in this work. The authors take full responsibility for the originality and accuracy of the manuscript.
Appendix G Reproducibility Statement
We are committed to ensuring the reproducibility of our work. To this end, we provide the following information:
Datasets: All experiments in this study are conducted on publicly available datasets (e.g., MS-COCO, Pokemon, Flickr), ensuring that others can directly access the data we used.
Models: Our experiments are based on open-source diffusion model architectures, and no proprietary or restricted-access models are employed.
Code and Implementation: The implementation of our method, including preprocessing, training, and evaluation scripts, will be released to the public upon publication. We also provide the code in the supplementary materials.
Hyperparameters and Settings: Detailed hyperparameter configurations (e.g., batch size, learning rate, training iterations, $s$, $r_{t}$) are explicitly described in the paper and supplemental materials.
Evaluation Metrics: We report standard and widely adopted evaluation metrics (e.g., ASR, AUC, TPR@1%FPR) to ensure comparability with prior work.